100ml brandy contains 40ml ethanol then find water mole fraction

Grade:12th pass

1 Answers

Arun
5 years ago
Volume of solute (ethanol): 40 ml, Volume of solvent (water): 60 ml∴ Volume of solution = 40+60 = 100 ml,...Concentration = Volume of solute / Volume of solution = 40 /100 × 100 = 40 %